OpenStudy (anonymous):

Check my answer! Factor completely. 3x^3 – 48x A. 3(x + 4)(x – 4) B. 3x(x – 16) C. 3x(x – 4)(x – 4) D. 3x(x – 4)(x + 4) My answer: B. or D.

OpenStudy (anonymous):

This is the question. :) @asnaseer

OpenStudy (asnaseer):

what have you tried?

OpenStudy (asnaseer):

you seem to be unsure of which answer is correct - I wanted to know how you got to those two choices?

DivineSolar (divinesolar):

Work your answer out here. Show us how you got the answer choices.

OpenStudy (anonymous):

Okay. Let me narrow down my answer choice.

OpenStudy (asnaseer):

it is best if you list each of your steps so that we can help spot where you may have made a mistake

OpenStudy (anonymous):

Okay. So, let's do D. first.

OpenStudy (anonymous):

\[3x(x - 4)(x + 4)\]After thinking, I decided to do the FOIL Method with (x - 4)(x + 4).\[(x - 4)(x +4)~becomes~x^2 - 4x - 4x + 16\]So, then I multiply by 3x.\[3x(x^2 - 16)\]\[(3x \times x^2) - (3x \times 16)\]\[(3x^3) - (48x)~becomes~3x^3 - 48.\]I believe D. is the correct answer.\(~~~\Huge\ddot \smile\)

OpenStudy (asnaseer):

yes it is :)

OpenStudy (anonymous):

Sorry for taking almost ten minutes, it took me some time to work out the LaTeX. >_<

OpenStudy (anonymous):

Really? :D :D :D Thank you for checking it. :D

DivineSolar (divinesolar):

Yes @BlossomCake

OpenStudy (asnaseer):

although you were probably expected to work from the equation given to you and arrive at one of the choices. You have worked in reverse by taking one of the choices and then showing it expands out to be the correct equation.

OpenStudy (anonymous):

Yes, sometimes I work better that way. :) It is also because I was unsure of how to work from the equation given to me.

OpenStudy (asnaseer):

I would have expected you to tackle this as follows:\[3x^3-48x=3(x^3-16x)=3x(x^2-16)=3x(x-4)(x+4)\]the last step involved using the difference of two squares, i.e.:\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (asnaseer):

NOTE: \(16=4^2\)

OpenStudy (anonymous):

Oh, yeah. Now that I look at it, that way seems simpler. Hm. :)

OpenStudy (asnaseer):

the more you practice these types of questions the better you will get at them - I am sure you'll be an expert in no time at all :)

OpenStudy (anonymous):

Well, I have one more problem but it is for a trinomial instead of a binomial... Maybe you can check that one, too. :)

OpenStudy (asnaseer):

sure - just post it as a new question in the usual way :)